OpenStudy (anonymous):

Write an equation of the line in point slope form, that passes through the two given points (-2, 15) and (9, -18)? (Not sure how to do this... help?)

OpenStudy (oaktree):

So do you know the equation for slope?

OpenStudy (anonymous):

Yes y2 - y1 over x2 - x1

OpenStudy (oaktree):

Right. So what do you get for slope?

OpenStudy (anonymous):

Just a minute...

OpenStudy (anonymous):

-33/11

OpenStudy (oaktree):

Right, now simplify that.

OpenStudy (anonymous):

-3

OpenStudy (oaktree):

Good. So now we have the equation y=-3x+b. Good so far?

OpenStudy (anonymous):

Yup

OpenStudy (oaktree):

Great. Now we choose a point and plug it in. Take the first one. We have 15=-3(2)+b. Now solve for b. What do you get?

OpenStudy (oaktree):

Correction: 15=-3(-2)+b. Left out a negative. Sorry.

OpenStudy (anonymous):

No problem... Do I multiply -2 by -3 first or divide -3 from both sides?

OpenStudy (oaktree):

Multiply.

OpenStudy (oaktree):

You should then get 15=6+b, and then subtract 6.

OpenStudy (anonymous):

Ok Thanks So it's 9 = b

OpenStudy (oaktree):

Right. So our solution is y=-3x+6.

OpenStudy (anonymous):

But it's in point slope form not slope intercept?

OpenStudy (oaktree):

Oh, sorry. So then we want to write it in point slope. Can you do that?

OpenStudy (anonymous):

I'm not sure how really.

OpenStudy (oaktree):

It's just\[y - 15 = -3(x+2)\]That's it.

OpenStudy (anonymous):

Ok thank you I get that!

OpenStudy (oaktree):

Great. Glad to help. Thanks for the medal!

OpenStudy (anonymous):

No problem, um I have another quick question.

OpenStudy (oaktree):

Fire away.

OpenStudy (anonymous):

So if we were using point slope form and we knew the slope was 3 and 2 points were (8, -8) how would we write that?

OpenStudy (anonymous):

Would it be y - 8 = 3 (x + 8) ?

OpenStudy (oaktree):

We would write \[y+8 = 3(x-8)\]You messed up your signs.

OpenStudy (anonymous):

Oops

OpenStudy (anonymous):

Thank you I meant to have it the other way around!

OpenStudy (oaktree):

Right.